# Sim-to-Real Transfer of Deep Reinforcement Learning Agents for Online Coverage Path Planning

### v2

@@ Conclusions
In this paper, we transfer state-of-the-art CPP policies to a semi-virtual environment in order to reduce the sim-to-real gap. We address the three challenges of sim-to-real transfer presented in the introduction. (1) The mismatch between simulated and real dynamics is reduced by a high inference frequency and fine-tuning the model in a real setting. (2) Non-Markovian dynamics are accounted for by using action observations, and incorporating realistic accelerations and delays into the simulation. (3) Computational delays are addressed by parallelizing the environment interactions and the model updates. In our experiments, we find that training the model in simulation assuming a first-order Markov process is sufficient, as long as the inference frequency is sufficiently high. Meanwhile, a higher-order policy can be further improved through fine-tuning, and can be deployed at a lower inference frequency, which lowers the computational requirements for a deployed system.
@@
Limitations and future work relate to accurate pose, simulated sensor data, and static environments with no moving obstacles. We discuss these further in the appendix.
Acknowledgements
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg (KAW) Foundation. The computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725, and by the Berzelius resource, provided by the KAW Foundation at the National Supercomputer Centre (NSC).
Appendix
This appendix contains the evaluation maps, extended results and analysis, as well as a discussion on limitations.
Evaluation Maps
Figure 5 illustrates the six evaluation maps used to measure the performance of the CPP policies.
Refer to caption Refer to caption Refer to caption
Map 1 Map 2 Map 3
Refer to caption Refer to caption Refer to caption
Map 4 Map 5 Map 6
Figure 5: The six evaluation maps used in our experiments.
Learned Paths
Figure 6 shows learned paths on four of the evaluation environments, and videos can be found online. [^1] Videos are available at: https://drive.google.com/drive/folders/1J0vpOBuRhCHOxsnAE1Vd1cZ6-qHPlvUh?usp=sharing
Optimal Training Step Size
In Table 4, we investigate the impact of the training step size in simulation, both for first-order and higher-order policies. In both cases, $500$ ms seems to be the best choice, considering $90\%$ and $99\%$ coverage times. Thus, we chose $500$ ms for the experiments in the main paper.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Learned paths showing the starting position (red triangle) and the end position (green square).
First order Higher order
$\Delta t$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
150 7.1 12.7 7.6 13.7
250 6.4 9.3 6.4 10.1
500 6.0 9.1 5.8 8.5
1000 6.1 9.6 6.4 9.1
Table 4: Training step size comparison ($\Delta t$ in ms). $T_{90}$ and $T_{99}$: Average time in minutes for reaching $90\%$ and $99\%$ coverage on the six evaluation maps in simulation.
@@ Acknowledgements
Table 5: First-order policy. Time in minutes for reaching $90\%$ and $99\%$ coverage. Env: Evaluation environment. Steps: Fine-tuning steps. $\Delta t$: Inference step size (ms). Note: Training step size is always 500 ms.
@@
Table 6: Higher-order policy. Time in minutes for reaching $90\%$ and $99\%$ coverage. Env: Evaluation environment. Steps: Fine-tuning steps. $\Delta t$: Inference step size (ms). Note: Training step size is always 500 ms.
If the step size is small, it may take longer to train in simulation since the computation time per step is the same. More steps are needed to complete the same number of episodes. In Table 4, we used the same number of training steps, which meant that the number of episodes were lower for the smaller step sizes. Meanwhile, with a lower step size, future rewards are more strongly discounted, and thus, long-term planning might be limited. It is possible that this could be remedied by tuning hyperparameters, such as the discount factor, $\gamma$. We simply used the default value of $\gamma=0.99$ for SAC (Haarnoja et al. 2018a). Finally, a lower step size requires a higher number of good consecutive actions to a achieve a good result, which might inhibit learning.
On the other hand, if the step size is too large, the short-term control is limited, as each action is applied over a longer time period. This requires the agent to be more precise in its actions, which makes the problem harder. A step size of $500$ ms seemed to strike a good balance.
Extended Coverage Results
@@
Tables 5 and 6 are extensions of Tables 2 and 3 in the main paper. They contain the $90\%$ and $99\%$ coverage times for each individual evaluation map, as well as the average times.
Average Speed
In Figure 7, we analyse the average speed of the agent during fine-tuning. The initial drop in performance is also evident here. In the initial fine-tuning phase, the agent attains a lower average speed, as a result of e.g. more collisions and many small turns, while it adjusts to the new environment dynamics. Over time, it learns to control the robot in a more efficient manner, allowing it to keep a higher average speed with less turns, which results in lower coverage times.
Learned Entropy
Figure 8 shows the learned entropy coefficient in the soft actor-critic framework (Haarnoja et al. 2018b). It increases initially, suggesting that the agent prioritizes exploration over exploitation during the distribution shift to the real world. After the initial peak, the entropy coefficient decreases, putting more emphasis on exploitation. However, it does not reach the same level as in simulation, i.e. at 00 fine-tuning steps. This either suggests that the real environment is more uncertain, or that the learning process has not entirely converged yet. It is likely a combination of the two.
Figure 7: Average speed during fine-tuning on evaluation maps 1 and 2. Dashed line: No fine-tuning.
Figure 8: The learned entropy coefficient in soft actor-critic (Haarnoja et al. 2018b) during fine-tuning.
Collision Frequency
The collision frequency on the six evaluation maps in the real setting was measured to be once every $6.2$ minutes, or once every $65.4$ meters, on average. However, the collisions were mainly from the side of the robot, and not head on. This is in line with the simulation results by Jonnarth, Zhao, and Felsberg (2024).
Additional Hyperparameters
In Table 7, we list additional hyperparameters used in our experiments. They are the same as used by Jonnarth, Zhao, and Felsberg (2024). Refer to the cited work for additional details on these parameters.
Parameter Description Value
Multi-scale maps
- grid resolution* (meters per pixel) $0.0375$
- map size** (number of grid cells) $32\times 32$
$m$ number of scales $4$
$s$ scale factor $4$
Reward parameters
$\lambda_{\mathrm{area}}$ maximum coverage reward $1$
$\lambda_{\mathrm{TV}}$ total variation reward scale $1$
$R_{\mathrm{coll}}$ collision penalty $-10$
$R_{\mathrm{const}}$ constant reward $-0.1$
Table 7: A list of additional hyperparameters used in our experiments. *Resolution for the finest scale, which is the same as the resolution for the full global map. **Note: this is not the size of the full environment, but rather the size of each scale in the multi-scale map representation.
Limitations
Accurate pose. The experiments are conducted with highly accurate pose estimates from an indoor positioning system, which may not be available, depending on the application. In such cases it needs to be estimated by other means, such as SLAM, which introduces higher levels of noise. Although the pose estimates in our experiments contain a low level of noise, Jonnarth, Zhao, and Felsberg (2024) show that adding high levels of localization and sensing noise does not impact performance to a significant extent. Thus, we believe that our conclusions likely extend to real environments with higher noise levels.
Sensing. The sim-to-real gap is mainly reduced with respect to robot kinematics and dynamics, while the sensor data remains simulated. This was done mainly to evaluate the sim-to-real transfer capability of the learning algorithm in a controlled manner, as stated in the main paper. Nevertheless, we believe that our conclusions likely extend to real sensors as well, for the same reason as above. Meanwhile, the sensing is omnidirectional in our experiments, which can be a strong assumption for some applications. However, the authors of the considered RL approach (Jonnarth, Zhao, and Felsberg 2024) demonstrated that it works even under non-omnidirectional sensing.
Dynamic obstacles. While dynamic obstacles are not explicitly modeled in the training framework, they can be accounted for by including, for instance, an object detector, to update the map based on detected moving obstacles. When an object is detected, the corresponding location in the map is replaced with non-free space in the obstacle map, and when the object is no longer detected, it is replaced with free space again. This was briefly tested in our experiments, with the following results. The robot did not collide with detected obstacles, but rather it planned around them. When the objects were moved away, and the robot explored the same region again, it planned its path through that region if necessary. In future work, we have to consider such experiments more systematically.

### v3

@@ VI Conclusions
Furthermore, we transfer the trained CPP policies to a semi-virtual environment in order to reduce the sim-to-real gap. We address the three challenges of sim-to-real transfer presented in the introduction. (1) The mismatch between simulated and real dynamics is reduced by a high inference frequency and fine-tuning the model in a real setting. (2) Non-Markovian dynamics are accounted for by using action observations, and incorporating realistic accelerations and delays into the simulation. (3) Computational delays are addressed by parallelizing the environment interactions and the model updates. In our experiments, we find that training the model in simulation assuming a first-order Markov process transfers well without fine-tuning, as long as the inference frequency is sufficiently high. Meanwhile, a higher-order policy can be further improved through fine-tuning, and can be deployed at a lower inference frequency, which lowers the computational requirements for a deployed system.
VII Limitations and Future Work
@@
Limitations and future work relate to accurate pose, simulated sensor data, and static environments with no moving obstacles, which we discuss here.
Accurate pose. The real-world experiments are conducted with highly accurate pose estimates from an indoor positioning system, which may not be available, depending on the application. In such cases it needs to be estimated by other means, such as SLAM, which introduces higher levels of noise. However, our simulation experiments show that even under high levels of localization and sensing noise, the performance of our method is not impacted to a significant extent. Thus, we believe that our conclusions likely extend to real environments with higher noise levels.
Sensing. The sim-to-real gap is mainly reduced with respect to robot kinematics and dynamics, while the sensor data remains simulated. This was done to evaluate the sim-to-real transfer capability of the learning algorithm in a controlled manner. Nevertheless, we believe that our conclusions likely extend to real sensors as well, for the same reason as above.
Dynamic obstacles. While dynamic obstacles are not explicitly modeled in the training framework, they can be accounted for by including, for instance, an object detector, to update the map based on detected moving obstacles. When an object is detected, the corresponding location in the obstacle map is replaced with non-free space, and when the object is no longer detected, it is replaced with free space again. This was briefly tested in our experiments, with the following results. The robot did not collide with detected obstacles, but rather it planned around them. When the objects were moved away, and the robot explored the same region again, it planned its path through that region if necessary. In future work, we have to consider such experiments more systematically.
Appendix A Agent Architecture Details
As proposed for soft actor-critic (SAC) [13], we use two Q-functions that are trained independently, together with accompanying target Q-functions whose weights are exponentially moving averages. No separate state-value network is used. All networks, including actor network, Q-functions, and target Q-functions share the same network architecture, which is either a multilayer perceptron (MLP), a naive convolutional neural network (CNN), or our proposed scale-grouped convolutional neural network (SGCNN). These are described in the following paragraphs. The MLP contains a total of 3.23.2M parameters, where most are part of the initial fully connected layer, which takes the entire observation vector as input. Meanwhile, the CNN-based architectures only contain 0.80.8M parameters, as the convolution layers process the maps with fewer parameters.
MLP. As mentioned in the main paper, the map and sensor feature extractors for MLP simply consist of flattening layers that restructure the inputs into vectors. The flattened multi-scale coverage maps, obstacle maps, frontier maps, and lidar detections are concatenated and fed to the fusion module. For the Q-functions, the predicted action is appended to the input of the fusion module. The fusion module consists of two fully connected layers with ReLU and 256 units each. A final linear prediction head is appended, which predicts the mean and standard deviation for sampling the action in the actor network, or the Q-value in the Q-functions.
Naive CNN. This architecture is identical to SGCNN, see below, except that it convolves all maps simultaneously in one go, without utilizing grouped convolutions.
SGCNN. Our proposed SGCNN architecture uses the same fusion module as MLP, but uses additional layers for the feature extractors, including convolutional layers for the image-like maps. Due to the simple nature of the distance readings for the lidar sensor, the lidar feature extractor only uses a single fully connected layer with ReLU. It has the same number of hidden neurons as there are lidar rays. The map feature extractor consists of a $2\times 2$ convolution with stride $2$ to reduce the spatial resolution, followed by three $3\times 3$ convolutions with stride 1 and without padding. We use $24$ total channels in the convolution layers. Note that the maps are grouped by scale, and convolved separately by their own set of filters. The map features are flattened into a vector and fed to a fully connected layer of $256$ units, which is the final output of the map feature extractor. ReLU is used as the activation function throughout.
Computational efficiency. As our goal is to perform fine-tuning and inference on edge hardware in real time, the computational efficiency of our model is highly important. In this regard, our SGCNN architecture is suitable, with only 0.80.8M parameters. This is lightweight compared to popular ImageNet computer vision models, such as MobileNet [53], ResNet [54], ConvNeXt [55], SwinTransformer [56, 57], and VisionTransformer [58]. In Table XI, we compare our model against these architectures in terms of number of parameters and inference time. The inference time for the ImageNet models were averaged over $1000$ forward passes on $64\times 64\times 3$-dimensional input images, as this corresponds to the same size as our $32\times 32\times 4\times 3$-dimensional input maps. However, the VisionTransformer expects a $224\times 224\times 3$-dimensional input, so this was used instead. As we can see from the table, our SGCNN architecture measures the lowest inference time with the fewest parameters.
TABLE XI: Model comparison in terms of number of parameters and inference time between our proposed SGCNN architecture and existing ImageNet computer vision models. The inference time is measured on two laptops; L1 (i5-520M CPU, no GPU) and L2 (i7-13700H CPU, 4060 GPU).
Model #params Time (L1) Time (L2)
MobileNet-V3-Small 2.5M 17 ms 8 ms
MobileNet-V3-Large 5.5M 25 ms 10 ms
ResNet-18 12M 31 ms 5 ms
ResNet-50 26M 77 ms 8 ms
ConvNeXt-Tiny 29M 82 ms 15 ms
Swin-V2-T 29M 175 ms 35 ms
ViT-Base-16 86M 1194 ms 116 ms
SGCNN (ours) 0.8M 5 ms 2 ms
Appendix B Choice of Hyperparameters
Multi-scale map parameters. We started by considering the size and resolution in the finest scale, such that sufficiently small details could be represented in the nearest vicinity. We concluded that a resolution of $0.0375$ meters per pixel was a good choice as it corresponds to $8$ pixels for a robot with a diameter of $30$ cm. Next, we chose a size of $32\times 32$ pixels, as this results in the finest scale spanning $1.2\times 1.2$ meters. Next, we chose the scale factor $s=4$, as this allows only a few maps to represent a relatively large region, while maintaining sufficient detail in the second finest scale. Finally, the training needs to be done with a fixed number of scales. The most practical way is to simply train with sufficiently many scales to cover the maximum size for a particular use case. Thus, we chose $m=4$ scales, which in total spans a $76.8\times 76.8$ m region, and can contain any of the considered training and evaluation environments. If the model is deployed in a small area, the larger scales would not contain any frontier points in the far distance, but the agent can still cover the area by utilizing the smaller scales. For larger use cases, increasing the size of the represented area by adding more scales is fairly cheap, as the computational cost is $\mathcal{O}(\log n)$.
Reward parameters. We started by fixing $\lambda_{\mathrm{area}}=1$ and tuned the learning rate accordingly. Due to the normalization of the area and TV rewards, they are given approximately the same weight for $\lambda_{\mathrm{area}}=1$ and $\lambda_{\mathrm{TV}}=1$. Thus, we used this as a starting point, which seemed to work well for the lawn mowing problem, while a lower weight for the TV rewards was advantageous for exploration. Here, we made a parameter sweep over $\lambda_{\mathrm{TV}}^{\mathrm{I}}$ and $\lambda_{\mathrm{TV}}^{\mathrm{G}}$, with values in $\{0.1,0.2,0.5,1,2\}$, and found $0.2$ to be the best choice for exploration. For the collision reward, $R_{\mathrm{coll}}$, we found 10-10 to work best from $\{0,-1,-10\}$. For the constant reward, $R_{\mathrm{const}}$, a value of 0.1-0.1 was chosen out of $\{0,-0.1,-1\}$.
Episode truncation. For large environments, we found it important not to truncate the episodes too early, as this would hinder learning. If the truncation parameter $\tau$ was set too low, the agent would not be forced to learn to cover an area completely, as the episode would simply truncate whenever the agent could not progress, without a large penalty. With the chosen value of $\tau=1000$, the agent would be greatly penalized by the constant reward $R_{\mathrm{const}}$ for not reaching the goal coverage, and would be forced to learn to cover the complete area in order to maximize the reward.
Appendix C Progressive Training
To improve the convergence in the early parts of training, we apply curriculum learning [59], which has shown to be effective in reinforcement learning [60]. We use simple maps and increase their difficulty progressively. To this end, we rank the fixed training maps by difficulty, and assign them into tiers, see Fig. 17. The maps in the lower tiers have smaller sizes and simpler obstacles. For the higher tiers, the map size is increased together with the complexity of the obstacles.
TABLE XII: First-order policy. Time in minutes for reaching $90\%$ and $99\%$ coverage. Env: Evaluation environment. Steps: Fine-tuning steps. Inference step size is in ms. Note: Training step size is always 500 ms.
@@ Appendix C Progressive Training
TABLE XIII: Higher-order policy. Time in minutes for reaching $90\%$ and $99\%$ coverage. Env: Evaluation environment. Steps: Fine-tuning steps. Inference step size is in ms. Note: Training step size is always 500 ms.
@@
TABLE XIV: The progressive training levels contain maps of increasingly higher tiers and goal coverage percentages. At the highest levels, generated maps with random floor plans and obstacles are also used.
Exploration Lawn mowing
Map Random Goal Map Random Goal
Level tiers maps coverage tiers maps coverage
[…]
2 1,2,4 $90\%$ 0,1 $90\%$
3 1,2,4 $95\%$ 0,1 $95\%$
4 1,2,4 $97\%$ 0–2 $95\%$
5 1,2,4 $99\%$ 0–2 $97\%$
6 1–4 $99\%$ 0–2 $99\%$
7 1–4 $99\%$ 0–3 $99\%$
8 1–5 $99\%$ 0–3 $99\%$
Furthermore, we define levels containing certain sets of maps and specific goal coverage percentages, see Table XIV. The agent starts at level 1, and progresses through the levels during training. To progress to the next level, the agent has to complete each map in the current level, by reaching the specified goal coverage. Note that randomly generated maps are also used in the higher levels, in which case the agent has to additionally complete a map with random floor plans, and a map with randomly placed circular obstacles to progress. Whenever random maps are used in a level, either a fixed map or a random map is chosen with a $50\%$ probability each at the start of every episode.
Appendix D Random Map Generation
Inspired by procedural environment generation for reinforcement learning [61], we use randomly generated maps during training to increase the variation in the training data and to improve generalization. We consider random floor plans and randomly placed circular obstacles. First, the side length of a square area is chosen uniformly at random, from the interval $[2.4,7.5]$ meters for mowing and $[9.6,15]$ meters for exploration. Subsequently, a random floor plan is created with a $70\%$ probability. Finally, with $70\%$ probability, a number of obstacles are placed such that they are far enough apart to guarantee that the agent can visit the entire free space, i.e. that they are not blocking the path between different parts of the free space. An empty map can be generated if neither a floor plan is created nor obstacles placed. In the following paragraphs, we describe the floor plan generation and obstacle placement in more detail.
Random floor plans. The random floor plans contain square rooms of equal size in a grid-like configuration, where neighboring rooms can be accessed through door openings. On occasion, a wall is removed completely or some openings are closed off to increase the variation further. First, floor plan parameters are chosen uniformly at random, such as the side length of the rooms from $[1.5,4.8]$ meters, wall thickness from $[0.075,0.3]$ meters, and door opening size from $[0.6,1.2]$ meters. Subsequently, each vertical and horizontal wall spanning the whole map either top-to-bottom or left-to-right is placed with a $90\%$ probability. After that, door openings are created at random positions between each neighboring room. Finally, one opening is closed off at random for either each top-to-bottom spanning vertical wall or each left-to-right spanning horizontal wall, not both. This is to ensure that each part of the free space can be reached by the agent.
Random circular obstacles. Circular obstacles with radius $0.25$ meters are randomly scattered across the map, where one obstacle is placed every four square meters on average. If the closest distance between a new obstacle and any previous obstacle or wall is less than $0.6$ meters, the new obstacle is removed to ensure that large enough gaps are left for the agent to navigate through and that it can reach every part of the free space.
Appendix E Extended Fine-tuning Coverage Results
@@
Tables XII and XIII are extensions of Tables VIII and IX in the main paper. They contain the $90\%$ and $99\%$ coverage times for each individual evaluation map, as well as the average times.
Tier 0 Refer to caption
Tier 1 Refer to caption Refer to caption Refer to caption Refer to caption
Tier 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Tier 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Tier 4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Tier 5 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: The fixed training maps used in simulation are grouped into tiers by difficulty, depending on their size and complexity of the obstacles.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 18: Training maps used during real-world fine-tuning.
(a) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(b) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(c) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(d) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(e) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 19: Evaluation maps used in simulation for (a) omnidirectional exploration, (b) non-omnidirectional exploration, and (c) lawn mowing. The maps in the last two rows were used additionally for ablations in (d) exploration and (e) lawn mowing.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 20: The evaluation maps used for the real-world experiments.
Appendix F Maps
In Fig. 17, 18, 19, and 20, we present the maps used in our experiments. Fig. 17 and 18 show the training maps used in simulation and the real-world, respectively. Fig. 19 and 20 show the evaluation maps used in simulation and the real-world, respectively. The maps are ordered from left to right in the same order as they appear in the respective tables in the main paper.
Appendix G Details on the Compared Methods
TSP-based solution with A*. For the TSP baseline we subdivide the environment into square grid cells of comparable size to the agent, and apply a TSP solver to find the shortest path to visit the center of each cell, which are the nodes. In order to guarantee that each cell is fully covered by a circular agent, the side length for each grid cell is set to $\sqrt{2}r$, where $r$ is the agent radius. Note that this leads to overlap and thus increases the time to reach full coverage. In the offline case, we compute the weight between each pair of nodes using the shortest path algorithm, A* [51]. However, due to the size of our environments, the computation time was infeasible. Thus, we implemented a supremum heuristic for distant nodes, where they would be assigned the largest possible path length instead of running A*. This improved the runtime considerably, while not affecting the path length to a noticeable extent. For the online case we applied the TSP solver on all visible nodes, executed the path while observing new nodes, and repeated until all nodes were covered. In this case, the TSP execution time on an Epyc 7742 64-core CPU was included in the coverage time, as the path would need to be replanned online in a realistic setting. However, the execution time was relatively small compared to the time to navigate the path. We also tried replanning in shorter intervals, but the gain in time due to a decreased path length was smaller than the increase in computation time.
Code repositories. Table XV lists the code implementations used for the methods under comparison. Note that for omnidirectional exploration, we evaluate our method under the same settings as in Explore-Bench, for which Xu et al. [6] report the performance of the compared methods. We report these results in Table III.
TABLE XV: Links to code implementations used for the methods under comparison.
Omnidirectional exploration
Official Explore-Bench implementation [6]:
https://github.com/efc-robot/Explore-Bench
Non-omnidirectional exploration
Official Frontier RL implementation [29]:
https://github.com/hanlinniu/turtlebot3_ddpg_collision_avoidance
Frontier RL implementation in simulation:
https://github.com/Peace1997/Voronoi_Based_Multi_Robot_Collaborate_Exploration_Unknow_Enviroment
Lawn mowing
Implementation of BSA [21]:
https://github.com/nobleo/full_coverage_path_planner